OpenStudy (anonymous):

What percent of 10 is 6? A. 60% B. 0.6% C. 1.67% D. 167%

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

a

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

6/10=.6

OpenStudy (anonymous):

6/10 = .6 = 60% A.

OpenStudy (anonymous):

\[\frac{ x }{ 100 }*10 = 6\] \[x=60%\]

OpenStudy (anonymous):

b

OpenStudy (anonymous):

b

OpenStudy (anonymous):

Best Answer 6/10% = 0.6% 6/10 = 3/5 = 60% a difference of 4% Not sure which one you want Hope this helped!!!